OpenStudy (anonymous):

If a sector in a circle of radius 10 has an area of 5 pie, what is the measure of the central angle that forms the sector? Please show work.

OpenStudy (anonymous):

area of circle: 100pi. sector area: 5pi 5/100=x/360 x=18

OpenStudy (dumbcow):

Area = pi*r^2*(angle/360) 5pi = 100pi*(angle/360) 1/20 = angle/360 angle = 18